# Optimal observables for (non-)equilibrium quantum metrology from the master equation

### v1

@@ Appendix B Expansion coefficients of the temperature SLD in the squeezed gaussian state
For completeness and reproducibility, we provide below the explicit expressions for the non-vanishing expansion coefficients of the temperature SLD in the squeezed Gaussian state, introduced in the main text as eqs. 47 and 48 and eq. 49. The expressions are also provided as Mathematica scripts in the supplementary material [24].
$c^{(x^{2})}_{T}=f(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{f_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{f_{D}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (60)
@@
$f_{D}$ $=\gamma\left[18\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(-72T\left\langle p^{2}\right\rangle+72T^{2}-4\right)+9T\right)\right.$ (62)
$\left.-18\gamma\left\langle\left\{x,p\right\}\right\rangle\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(-9\left\langle p^{2}\right\rangle\left(4T\left\langle x^{2}\right\rangle+1\right)+\left(72T^{2}-4\right)\left\langle x^{2}\right\rangle+18T\right)-9\right)+9T-4\left\langle x^{2}\right\rangle\right)\right.$
$\left.-9\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(9\left(\gamma^{2}+\left\langle p^{2}\right\rangle^{2}\right)+4\left\langle x^{2}\right\rangle\left(2T-3\left\langle p^{2}\right\rangle\right)\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(9\left(9\gamma^{2}+2\right)+4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle\left(9\gamma^{2}+4\left\langle x^{2}\right\rangle\left(4T-3\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-2T\right)+1\right)+18T-4\left\langle x^{2}\right\rangle\right)\right)\right.$
@@
$c^{(p^{2})}_{T}=g(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{g_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{g_{D}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (63)
$g_{N}$ $=9\left[-9\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(8T-4\left\langle p^{2}\right\rangle\right)+1\right)-72\gamma\left\langle x^{2}\right\rangle\left\langle\left\{x,p\right\}\right\rangle\left(T-\left\langle p^{2}\right\rangle\right)\left(4T\left\langle x^{2}\right\rangle+1\right)\right.$ (64)
$\left.-\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(T-\left\langle p^{2}\right\rangle\right)\left(8\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle-2\left\langle x^{2}\right\rangle\right)+9\right)-\left(16\left\langle x^{2}\right\rangle^{2}+9\right)\left(T-\left\langle p^{2}\right\rangle\right)\left(4\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+1\right)-9\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{5}\right]$
$g_{D}$ $=2\gamma\left[18\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(-72T\left\langle p^{2}\right\rangle+72T^{2}-4\right)+9T\right)\right.$
$\left.-18\gamma\left\langle\left\{x,p\right\}\right\rangle\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(-9\left\langle p^{2}\right\rangle\left(4T\left\langle x^{2}\right\rangle+1\right)+\left(72T^{2}-4\right)\left\langle x^{2}\right\rangle+18T\right)-9\right)+9T-4\left\langle x^{2}\right\rangle\right)\right.$
$\left.-9\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(9\left(\gamma^{2}+\left\langle p^{2}\right\rangle^{2}\right)+4\left\langle x^{2}\right\rangle\left(2T-3\left\langle p^{2}\right\rangle\right)\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(9\left(9\gamma^{2}+2\right)+4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle\left(9\gamma^{2}+4\left\langle x^{2}\right\rangle\left(4T-3\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-2T\right)+1\right)+18T-4\left\langle x^{2}\right\rangle\right)\right)\right.$
$c^{(\{x,p\})}_{T}=h(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{h_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{h_{D}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (65)
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(4\left\langle x^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-T\right)-9\gamma^{2}\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle\left(T-\left\langle p^{2}\right\rangle\right)\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle+4\left\langle x^{2}\right\rangle\right)+9\right)+4\left(1-9\gamma^{2}\right)\left\langle x^{2}\right\rangle\right)+9\gamma\left\langle p^{2}\right\rangle\left\langle\left\{x,p\right\}\right\rangle{}^{4}\right]$
$h_{D}$ $=\gamma\left[18\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(-72T\left\langle p^{2}\right\rangle+72T^{2}-4\right)+9T\right)\right.$ (67)
$\left.-18\gamma\left\langle\left\{x,p\right\}\right\rangle\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(-9\left\langle p^{2}\right\rangle\left(4T\left\langle x^{2}\right\rangle+1\right)+\left(72T^{2}-4\right)\left\langle x^{2}\right\rangle+18T\right)-9\right)+9T-4\left\langle x^{2}\right\rangle\right)\right.$
$\left.-9\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(9\left(\gamma^{2}+\left\langle p^{2}\right\rangle^{2}\right)+4\left\langle x^{2}\right\rangle\left(2T-3\left\langle p^{2}\right\rangle\right)\right)\right.$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(9\left(9\gamma^{2}+2\right)+4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle\left(9\gamma^{2}+4\left\langle x^{2}\right\rangle\left(4T-3\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-2T\right)+1\right)+18T-4\left\langle x^{2}\right\rangle\right)\right)\right.$

### v2

@@ Appendix B Expansion coefficients of the temperature SLD in the squeezed gaussian state
For completeness and reproducibility, we provide below the explicit expressions for the non-vanishing expansion coefficients of the temperature SLD in the squeezed Gaussian state, introduced in the main text as eqs. 46 and 47 and eq. 48. The expressions are also provided as Mathematica scripts in the supplementary material [24].
$c^{(x^{2})}_{T}$ $=f(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{f_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{\Delta(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (59)
$c^{(p^{2})}_{T}$ $=g(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{g_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{\Delta(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (60)
$c^{(\{x,p\})}_{T}$ $=h(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)=\frac{h_{N}(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}{\Delta(\langle x^{2}\rangle,\langle p^{2}\rangle,\langle\{x,p\}\rangle)}$ (61)
@@
$g_{N}$ $=\tfrac{9}{2}\left[-9\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(8T-4\left\langle p^{2}\right\rangle\right)+1\right)\right.$ (63)
$\left.-72\gamma\left\langle x^{2}\right\rangle\left\langle\left\{x,p\right\}\right\rangle\left(T-\left\langle p^{2}\right\rangle\right)\left(4T\left\langle x^{2}\right\rangle+1\right)\right.$
$\left.-\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(T-\left\langle p^{2}\right\rangle\right)\left(8\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle-2\left\langle x^{2}\right\rangle\right)+9\right)\right.$
$\left.-\left(16\left\langle x^{2}\right\rangle^{2}+9\right)\left(T-\left\langle p^{2}\right\rangle\right)\left(4\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+1\right)-9\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{5}\right]$
+$\left.+\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(4\left\langle x^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-T\right)-9\gamma^{2}\right)+\left\langle\left\{x,p\right\}\right\rangle\left(T-\left\langle p^{2}\right\rangle\right)\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle+4\left\langle x^{2}\right\rangle\right)\right.\right.\right.$
$\left.\left.\left.+9\right)+4\left(1-9\gamma^{2}\right)\left\langle x^{2}\right\rangle\right)+9\gamma\left\langle p^{2}\right\rangle\left\langle\left\{x,p\right\}\right\rangle{}^{4}\right]$
$\Delta$ $=\gamma\left[18\gamma\left\langle\left\{x,p\right\}\right\rangle{}^{3}\left(\left\langle x^{2}\right\rangle\left(-72T\left\langle p^{2}\right\rangle+72T^{2}-4\right)+9T\right)\right.$ (65)
$\left.-18\gamma\left\langle\left\{x,p\right\}\right\rangle\left(\left\langle p^{2}\right\rangle\left(4\left\langle x^{2}\right\rangle\left(-9\left\langle p^{2}\right\rangle\left(4T\left\langle x^{2}\right\rangle+1\right)+\left(72T^{2}-4\right)\left\langle x^{2}\right\rangle+18T\right)-9\right)\right.\right.$
$\left.\left.+9T-4\left\langle x^{2}\right\rangle\right)-9\left\langle\left\{x,p\right\}\right\rangle{}^{4}\left(9\left(\gamma^{2}+\left\langle p^{2}\right\rangle^{2}\right)+4\left\langle x^{2}\right\rangle\left(2T-3\left\langle p^{2}\right\rangle\right)\right)+\left\langle\left\{x,p\right\}\right\rangle{}^{2}\left(9\left(9\gamma^{2}+2\right)\right.\right.$
$\left.\left.+4\left\langle x^{2}\right\rangle\left(9\left\langle p^{2}\right\rangle\left(9\gamma^{2}+4\left\langle x^{2}\right\rangle\left(4T-3\left\langle p^{2}\right\rangle\right)+9\left\langle p^{2}\right\rangle\left(\left\langle p^{2}\right\rangle-2T\right)+1\right)+18T-4\left\langle x^{2}\right\rangle\right)\right)\right.$